OpenStudy (anonymous):

One more question? Easy equation, will give medals. 6n + 11 = -7 + 3n. Solve for n

OpenStudy (anonymous):

Please help?

OpenStudy (anonymous):

Thanks Kellie :)

OpenStudy (anonymous):

6n + 11 = -7 + 3n -- subtract 3n from both sides 6n - 3n + 11 = -7 -- subtract 11 from both sides 6n - 3n = -7 - 11 -- combine like terms 3n = - 18 -- divide both sides by 3 n = -6 check.. 6(-6) + 11 = -7 + 3(-6) -36 + 11 = -7 - 18 -25 = -25 (correct) n = -6

OpenStudy (anonymous):

do you have any questions ?

OpenStudy (anonymous):

Thats the equation I have, lol. I just didn't know if it was right! :D Thnaks Kells, heres your medal

OpenStudy (anonymous):

glad to help :)

OpenStudy (anonymous):

and thank you :)

OpenStudy (anonymous):

Can I ask another question?

OpenStudy (anonymous):

What is the first step in solving the equation Eighty-eight equals thirteen plus five x

OpenStudy (anonymous):

88= 13 + 5x

OpenStudy (anonymous):

88 = 13 + 5x --- subtract 13 from both sides (first step)

OpenStudy (anonymous):

n=-6

OpenStudy (anonymous):

Thanks @kelliegirl33 And thanks for the answer I never asked @TazMann

OpenStudy (anonymous):

How do you know thats the answer..?

OpenStudy (anonymous):

no problem :)

OpenStudy (anonymous):

yea sorry bout that didnt see the other question

OpenStudy (anonymous):

It's okay.. One more question for luck? Lol 3x - 6 + 15

OpenStudy (anonymous):

3x - 6 + 15 ---all we can do is combine like terms because there is no equal sign. 3x + 9 (I do not know if you just want this.. this can be factored, but I do not know if your teacher wants you to do this.. 3x + 9 3(x + 3) --- this is also a possible answer if your teacher wants you to factor

OpenStudy (anonymous):

or did you write the problem wrong and there is supposed to be an equal sign ?

OpenStudy (anonymous):

Hm, Idk. I think thats the answer.

OpenStudy (anonymous):

Wait

OpenStudy (anonymous):

wow

OpenStudy (anonymous):

3x - 6 = 15 XDDD

OpenStudy (anonymous):

Shahtup Taz :P

OpenStudy (anonymous):

my apologies

OpenStudy (anonymous):

3x - 6 = 15 3x = 15 + 6 3x = 21 x = 7

OpenStudy (anonymous):

What are you English? @TazMann

OpenStudy (anonymous):

DANKA Thanks in german THANKS :D

OpenStudy (anonymous):

nope i love math

OpenStudy (anonymous):

Lol, "my apologies" :P XD

OpenStudy (anonymous):

25 - 46 = -6b + 47

OpenStudy (anonymous):

nice haha

OpenStudy (anonymous):

:P Hewp

OpenStudy (anonymous):

25 - 46 = -6b + 47 ---Is this the problem you need help on ? Is it written correctly ?

OpenStudy (anonymous):

It tiz

OpenStudy (anonymous):

25 - 46 = -6b + 47 combine like terms -21 = -6b + 47 -- subtract 47 from both sides -47 - 21 = -6b -68 = -6b -- divide by -6 -68/-6 = b 68/6 which reduces to 34/3 or if you need a mixed number, it is 11 1/3